VEGAS Team Meeting

GB-137-hub
192.33.117.12##7144
434-817-6443# (telephone)

Attendees:

Adam, Marty, John, Richard, Joe, Paul, Jason, Mark, Dave, Ray, Anish

Agenda:

  1. Status of internal RFI and OGP solution
  2. Mode 10-19 hardware fix
  3. Baseline ringing (high frequency noise) in VEGAS data
  4. Integration time MR
  5. Top Hat
  6. Test Time
  7. Other

Meeting notes:

  1. Status of internal RFI and OGP solution
    • Waiting for UCB
    • Solution:
      • Flagging for LBW modes - locations not as simply located as for HBW modes
      • Implementation details needed [Adam]
        • Joe can assist w/ details
      • MR 1 in preparation - solution
      • MR 2 in preparation - SDFITS changes
    • We can suspend weekly tests
  2. Mode 10-19 hardware fix
    • Testing Tuesday
      • High channel mode fix (v.3)
      • Mode 4-19 Fixed bug in switching interpretation (data vs. phase state)
  3. Baseline ringing (high frequency noise) in VEGAS data
    • Status of more stable and insulated cabling
      • Electronics is fabrication a couple higher quality cables to test
    • (hopefully) testing the S/N ration range of VEGAS with attenuation on Thursday
      • Determine sufficient S/N ratio to support use of 20dB attenuators to improve ripples
  4. Integration time MR
    • Ray has drafted
    • Joe is checking
  5. Top Hat -
    • change to abort scan upon occurrence
    • any update/occurrence? [No]
  6. Test Time
    • Tuesday
      • Mode 10-19 hardware fix
    • Thursday
      • S/N ratio tests
  7. Other

-- MartyBloss - 2014-10-06
Topic revision: r2 - 2014-10-06, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback